OpenStudy (anonymous):

Is x = 1 a solution of the equation 2 - 8x = -6?

OpenStudy (austinl):

Plug in 1 for x in the equation. 2-8(1)=-6 2-8=-6 -6=-6 Therefore it is a valid solution.

OpenStudy (anonymous):

so its yes ?

OpenStudy (austinl):

Yep